1 / 4

LHC Operational Updates: Injection, Steering, and Stability

Stay updated with the latest LHC operational activities including injections, beam steering adjustments, and achieving stable beams. Follow the progress and improvements for enhanced performance.

heaton
Download Presentation

LHC Operational Updates: Injection, Steering, and Stability

An Image/Link below is provided (as is) to download presentation Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. Content is provided to you AS IS for your information and personal use only. Download presentation by click this link. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server. During download, if you can't get a presentation, the file might be deleted by the publisher.

E N D

Presentation Transcript


  1. Last 24 hours

  2. Saturday 20th October • 11:20 lost fill 3201 to spurious opening extraction switch of RCS.A12B1 (as Friday) • 123 pb-1 in 10h18m • Access for MPE piquet – switch replacement • Y.Bastian (QPS) joins Vincent Froidbisefor the intervention on RCS.A12.B1 • 13:30 Access to PM25 finished. Closing LHC... • 14:00 Ready for injection • SPS working on beam quality leading into… • Main Power Supply problem in the Booster LHC status

  3. Saturday 20th October • 17:00 Injection • Some steering required • QC latched on injection oscillation. steering on both beams correcting the lines but what improved the situation was to increase the SPS scraping. could inject but from time to time the IQC was latching… • 19:30 Start ramp • Without high bandwidth ADT • 20:20 Stable beams Fill 3202 • Initial luminosity: 7.02e33 cm-2s-1 • ALICE ~5e30 cm-2s-1 - will tweak up next week LHC status

  4. Sunday 21st October • 08:03 OP dump of 3203 • 172 pb-1 in 11h44m • 08:55 Re-injecting for physics LHC status

More Related